OpenStudy (geometry_hater):

Use quadratic formula to solve this equation 4x squared = 3x + 5

hartnn (hartnn):

\(4x^2=3x+5\) so\(4x^2-3x-5=0\) Compare this ^^ quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two rots of x are: \(x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\)